Integrate with respect to x :
∫10x^2dx
a. 4 1/3
b. 1
c. 2
d. 12
1
Expert's answer
2014-09-12T11:25:45-0400
You missed limits of integration, because statement concerns indefinite integral, answers concern the definite integral.
Calculation of indefinite integral ∫10x2dx=10∫x2dx=10x3/3+C, where C is an arbitrary real constant. If you integrate x2 from 1 to 0, then answer is (03/3-13/3)=-1/3. If you integrate x2 from 0 to 1, then answer is (13/3-03/3)=1/3.